OpenStudy (anonymous):

find the equation of the path described by the following vector. u=t^2+4/(t+2)^2 + 4t/(t+2)^2

OpenStudy (amistre64):

thats a mean looking vector

OpenStudy (anonymous):

lol yeah

OpenStudy (amistre64):

the left side is usually the 'x' component; which leaves the other to be the 'y' component

OpenStudy (amistre64):

they have the same bottoms on i see that

OpenStudy (anonymous):

it gives a hint to factorise but im not sure how to use it

OpenStudy (amistre64):

t^2 +4 + 4t (t+2)(t+2) ------------ = --------- = 1 (t+2) (t+2) (t+2)(t+2)

OpenStudy (amistre64):

y = 1 with a hole at -2 i think

OpenStudy (amistre64):

x+y = 1 y = -x then

OpenStudy (anonymous):

lol the answers is y=1-x

OpenStudy (amistre64):

...... or that lol

OpenStudy (anonymous):

wait yeah thats right

OpenStudy (anonymous):

didnt see the second answer lol

OpenStudy (amistre64):

it is? como es posible?

OpenStudy (anonymous):

lol thanks heaps ur a genius